# New phenomena arising from ${\mathcal{L}}$-invariants of modular forms

### v1

@@ 2.3. Central values and quadratic twists
Thus $L_{\infty}(f,j)\in\overline{{{\mathbb{Q}}}}$, which is viewed in ¯$\overline{{{\mathbb{Q}}}}_{p}$ via $\iota$. It is called an algebraic special value, the most important of which is the central value at $j=k/2$. The following exceptional zero formula expresses the ${\mathcal{L}}$-invariant of an eigenform in terms of the central values of its $p$-adic $L$-function and complex $L$-series.
@@
The obstructions to calculating ${\mathcal{L}}$-invariants using (EZ) are that the complex $L$-value could vanish and $a_{p}(f)$ could be negative (i.e. equal to $-p^{\frac{k}{2}-1}$). However, by (1), we can always simultaneously twist ourselves out of these two situations. Indeed, if $\chi$ is a quadratic Dirichlet character unramified at $p$, then $f_{\chi}$ is also an eigenform of level $\Gamma_{0}$, except possibly with a different tame level and $a_{p}(f_{\chi})=\chi(p)a_{p}(f)$. The following proposition guarantees the existence of a twist of $f$ whose $p$-th Fourier coefficient is positive and whose central $L$-value is non-zero.
@@
We first seek a twist of $f$ having $+1$ as the sign of its functional equation and having $+p^{k/2-1}$ as its $U_{p}$-eigenvalue. Said another way, we need a $D$ co-prime to $N$ such that
[…]
@@
In practice, to calculate ${\mathcal{L}}$-invariants, we first calculate the modular symbols $\lambda_{f}$ for $\Gamma_{0}$-eigenforms $f$. This is a rather expensive step. Luckily, if we ever need to apply Proposition 2.4, we do not need to re-do an expensive calculation. Indeed, the modular symbol for $f_{\chi}$ can be readily computed from the symbol for $f$, by following formula (see [56, (8.5)]): if $\chi$ has conductor $D$, then
@@ 3. Estimating the $p$-adic derivative $L_{p}^{\prime}(f,k/2)$
In this section, we construct a sequence $\{L_{n}\}$ of explicit $p$-adic approximations for $L_{p}^{\prime}(f,k/2)$. See also [69, Proposition 3.1], which treats the ordinary ($k=2$) case.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
with $g(x)$ a locally analytic function on ${\mathbb{Z}}_{p}^{\times}$. To understand how to approximate $L_{p}^{\prime}(f,k/2)$, let’s recall how one extends $\mu_{f}$ from a functional on the space of locally polynomial functions of degree at most $k-2$ (as in (3)) to a distribution on all locally analytic functions.
To this end, assume that $n$ is large enough so that $g(x)$ is analytic on $a+p^{n}{\mathbb{Z}}_{p}$ for all $a$. Then decompose ${\mathbb{Z}}_{p}^{\times}$ as a disjoint union of balls of the form $a_{i}+p^{n}{\mathbb{Z}}_{p}$ with each $a_{i}\in{\mathbb{Z}}$ (depending on $n$). Write $\operatorname{TS}_{a_{i}}(g)$ for the Taylor series expansion of $g$ around $x=a_{i}$ truncated to terms of degree at most $k-2$. One defines
[…]
@@
which by Lemma 3.2 and the above discussion converges to $L_{p}^{\prime}(f,k/2)$ as $n\to\infty$. Since $L_{n}$ is defined by values of $\mu_{f}$ on polynomials of degree at most $k-2$, we see from Section 2.2 that $L_{n}$ is a linear combination of period of integrals of $f$, and thus each $L_{n}$ can be computed from the modular symbol attached to $f$.
@@
[…]
@@
[…]
@@ 4. Summary of algorithm to calculate ${\mathcal{L}}$-invariants
    The eigenvalues of $U_{p}$ acting on $A$ must all be of the form ±$\pm p^{\frac{k}{2}-1}$ with a fixed sign. We calculate the $p$-th Hecke polynomial and use it to determine the sign of $a_{p}=a_{p}(A)$. Since $A$ is new of level $\Gamma_{0}$, the sign $w=w(A)$ of the functional equation can extracted from the sign of the eigenvalue for the Atkin–Lehner operator $w_{Np}$ on $A$. Having determined these two signs, we find a quadratic discriminant $D$ as in Proposition 2.4. Write $\chi=\chi_{D}$ for the quadratic character of conductor $D$.
@@
    is equal to $v_{\mathfrak{p}}({\mathcal{L}})$ for $[K_{\mathfrak{p}}:{\mathbb{Q}}_{p}]$-many eigenforms occurring in $A$. So, we store this value, with multiplicity $[K_{\mathfrak{p}}:{\mathbb{Q}}_{p}]$. If (8) fails for some $\mathfrak{p}$, then return to step (iii)(b) to compute at a deeper level, eventually replacing $L_{1}$ by $L_{2}$, and so on.
@@ 5.1. Initial plots in tame level one
  • The size of a scatter point represents the datum’s multiplicity. For instance if $v_{p}({\mathcal{L}})=-8$ occurs twice in weight $k$ and $v_{p}({\mathcal{L}})=-10$ occurs four times, then the scatter point at $(k,-10)$ is a magnitude larger than the scatter point at $(k,-8)$.
@@
Figure 5. $3$-adic slopes of ${\mathcal{L}}$-invariant slopes in level $\Gamma_{0}(3)$ with respect to weights $6\leq k\leq 508$. (Total data: 10752 eigenforms.)
Figure 6. $5$-adic slopes of ${\mathcal{L}}$-invariant slopes in level $\Gamma_{0}(5)$ with respect to weights $10\leq k\leq 472$. (Total data: 18560 eigenforms.)
Figure 7. $7$-adic slopes of ${\mathcal{L}}$-invariant slopes in level $\Gamma_{0}(7)$ with respect to weights $4\leq k\leq 290$. (Total data: 10511 eigenforms.)
@@ 5.3. The sign of $a_{p}$
These split as evenly as they could among the two signs: the $-9$ occurs once for each sign and the $-4$ occurs once for each sign. The $-2$ happens to occur for an eigenform with $a_{3}<0$. For a second example, in weight $k=28$ and level $\Gamma_{0}(3)$, there are an even number of eigenforms and their ${\mathcal{L}}$-invariants have slopes
@@
Both plots support our two principles above. For instance, in level $\Gamma_{0}(2)$ the most negative sign dependent ${\mathcal{L}}$-invariant slope we found was $-7$. This is extremely close to zero, once we note the most negative ${\mathcal{L}}$-invariant slope over the same data range is less than $-120$ (see Figure 3). It seems plausible to expect sign dependent slopes $v$ to satisfy $v\geq-c\log(k)$ where $c$ is a positive constant. (Depending on what? Perhaps just $p$.) It also seems reasonable that the total number of sign dependent slopes in weights $k\leq T$ is $O(T\log(T))$, which is a magnitude smaller than roughly $T^{2}$-many eigenforms of weight $k\leq T$.
@@
$p$ $Np$ weights # data up $k_{\mathrm{mid}}$ sign dependent up to $k_{\mathrm{mid}}$ total data % sign dependent
$2$ $2$ $391$ $3434$ $4.14$ $13002$ $2.37$
$2$ $6$ $165$ $1291$ $9.37$ $4603$ $5.02$
$2$ $10$ $136$ $1680$ $7.5$ $6264$ $3.64$
$2$ $14$ $126$ $2245$ $8.06$ $7947$ $3.49$
$2$ $22$ $103$ $2340$ $4.79$ $9013$ $2.34$
$3$ $3$ $252$ $2730$ $4.25$ $10752$ $2.18$
$3$ $6$ $164$ $1291$ $6.89$ $4546$ $4.4$
$3$ $15$ $72$ $962$ $12.47$ $3594$ $6.12$
$3$ $21$ $70$ $1296$ $9.57$ $5040$ $4.88$
$3$ $33$ $42$ $735$ $14.83$ $2941$ $7.17$
$5$ $5$ $243$ $5371$ $3.18$ $20341$ $1.61$
$5$ $10$ $135$ $1680$ $5.24$ $6173$ $2.64$
$5$ $15$ $72$ $962$ $8.94$ $3594$ $5.29$
$5$ $35$ $41$ $883$ $10.31$ $3363$ $6.87$
$5$ $70$ $20$ $280$ $9.29$ $960$ $5.62$
$7$ $7$ $144$ $2664$ $3.38$ $10512$ $1.77$
$7$ $14$ $90$ $1301$ $7.76$ $4085$ $4.38$
$7$ $21$ $70$ $1296$ $6.64$ $5040$ $3.37$
$7$ $35$ $41$ $883$ $10.76$ $3363$ $5.68$
$7$ $70$ $19$ $280$ $10.0$ $874$ $6.41$
@@ 5.4. Galois representations
Figure 13 offers a second plot to visualize how $\overline{\rho}$-filtered data sits within an unfiltered dataset. Namely, let $\overline{\rho}$ be the mod $11$ representation associated with the elliptic curve $X_{0}(11)$. Then we consider all eigenforms $f$ for which $\overline{\rho}_{f}\simeq\overline{\rho}\otimes\omega^{j}$ for some $j=0,1,\dotsc,10$. We call these eigenforms “twists of $X_{0}(11)$ modulo $11$”. Each such $\overline{\rho}_{f}$ is surjective, which makes the context slightly different than the Eisenstein examples above.
@@ 6.1. ${\mathcal{L}}$-invariants and local Galois representations
Fix $k\geq 2$, a sign $\pm$, and ¯${\mathcal{L}}\in\overline{{{\mathbb{Q}}}}_{p}$. We write $D=\overline{{{\mathbb{Q}}}}_{p}e_{1}\oplus\overline{{{\mathbb{Q}}}}_{p}e_{2}$ and define ¯$\overline{{{\mathbb{Q}}}}_{p}$-linear operators $\varphi$ and $N$ on $D$ by
$\varphi$ =(±$=\begin{pmatrix}\pm p^{\frac{k}{2}}&0\\ 0&\pm p^{\frac{k}{2}-1}\end{pmatrix}$
@@
In this way, $D$ is a two-dimensional ¯$\overline{{{\mathbb{Q}}}}_{p}$-linear weakly-admissible filtered $(\varphi,N)$-module in the sense of Fontaine’s $p$-adic Hodge theory of local Galois representations [39]. A theorem of Colmez and Fontaine ([33, Théormème A]) implies there exists a two-dimensional local representation
@@
that is semi-stable and non-crystalline, since $N\neq 0$ on $D$, for which $D_{\mathrm{st}}(r)=D$. The Hodge–Tate weights of $r$ equal $\{0,k-1\}$ and $\det(r)=\omega^{k-1}$. We write $r_{k,{\mathcal{L}}}^{\pm}$ for this representation, or we write $V_{k,{\mathcal{L}}}^{\pm}$ when it is more convenient to reference the vector space upon which the Galois groups acts. Any two-dimensional, semi-stable, and non-crystalline, ¯$\overline{{{\mathbb{Q}}}}_{p}$-linear $\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})$-representation with distinct Hodge–Tate weights is isomorphic to a twist of some such $r_{k,{\mathcal{L}}}^{\pm}$. The representations $r_{k,{\mathcal{L}}}^{\pm}$ are irreducible if and only if $k>2$. See [17, Exemple 3.1.2.2] for a similar summary and references.
@@
To avoid complications below, it is convenient here to include a definition of $r_{k,{\mathcal{L}}}^{\pm}$ for ${\mathcal{L}}=\infty$. Namely, $r=r_{k,\infty}^{\pm}$ is the unique two-dimensional, irreducible, crystalline representation of $\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})$ such that $D_{\mathrm{crys}}(r)=\overline{{{\mathbb{Q}}}}_{p}e_{1}\oplus\overline{{{\mathbb{Q}}}}_{p}e_{2}$ is the weakly-admissible filtered $(\varphi,N)$-module, with $N=0$, and for which

### v2

@@ 2.3. Central values and quadratic twists
Thus $L_{\infty}(f,j)\in\overline{{{\mathbb{Q}}}}$, which is viewed in $\overline{{{\mathbb{Q}}}}_{p}$ via $\iota$. It is called an algebraic special value, the most important of which is the central value at $j=k/2$. The following exceptional zero formula expresses the $\mathcal{L}$}-invariant of an eigenform in terms of the central values of its $p$-adic $L$-function and complex $L$-series.
@@
The obstructions to calculating $\mathcal{L}$}-invariants using (EZ) are that the complex $L$-value could vanish and $a_{p}(f)$ could be negative (i.e. equal to $p^{\frac{k}{2}-1}$). However, by (1), we can always simultaneously twist ourselves out of these two situations. Indeed, if $\chi$ is a quadratic Dirichlet character unramified at $p$, then $f_{\chi}$ is also an eigenform of level $\Gamma_{0}$, except possibly with a different tame level and $a_{p}(f_{\chi})=\chi(p)a_{p}(f)$. The following proposition guarantees the existence of a twist of $f$ whose $p$-th Fourier coefficient is positive and whose central $L$-value is non-zero.
@@
We first seek a twist of $f$ having +1+1 as the sign of its functional equation and having $p^{k/2-1}$ as its $U_{p}$-eigenvalue. Said another way, we need a $D$ co-prime to $N$ such that
(5) $\chi_{D}(-Np)=w_{f}\text{~~and~~}\chi_{D}(p)=\text{sign~of~}a_{p}(f).$
@@
In practice, to calculate $\mathcal{L}$}-invariants, we first calculate the modular symbols $\lambda_{f}$ for $\Gamma_{0}$-eigenforms $f$. This is a rather expensive step. Luckily, if we ever need to apply Proposition 2.4, we do not need to re-do an expensive calculation. Indeed, the modular symbol for $f_{\chi}$ can be readily computed from the symbol for $f$, by following formula (see [57, (8.5)]): if $\chi$ has conductor $D$, then
@@ 3. Estimating the $p$-adic derivative $L_{p}^{\prime}(f,k/2)$
In this section, we construct a sequence $\{L_{n}\}$ of explicit $p$-adic approximations for $L_{p}^{\prime}(f,k/2)$. See also [70, Proposition 3.1], which treats the ordinary ($k=2$) case.
@@
$L_{p}^{\prime}(f,k/2)=\int_{{\mathbb{Z}}_{p}^{\times}}x^{\frac{k}{2}-1}\cdot\log_{p}\langle x\rangle~\!d\mu_{f}(x).$
@@
$L_{p}(f,\omega^{\frac{k-2}{2}},s)$ $=\int_{{\mathbb{Z}}_{p}^{\times}}\omega(x)^{\frac{k}{2}-1}\langle x\rangle^{s-1}~\!d\mu_{f}(x)$
$=\int_{{\mathbb{Z}}_{p}^{\times}}x^{\frac{k}{2}-1}\langle x\rangle^{s-k/2}~\!d\mu_{f}(x)$
$=\int_{{\mathbb{Z}}_{p}^{\times}}x^{\frac{k}{2}-1}\exp((s-k/2)\log_{p}\langle x\rangle)~\!d\mu_{f}(x)$
@@
$=\sum_{n=0}^{\infty}\frac{1}{n!}\left(\int_{{\mathbb{Z}}_{p}^{\times}}x^{\frac{k}{2}-1}(\log_{p}\langle x\rangle)^{n}~\!d\mu_{f}(x)\right)(s-k/2)^{n}.$
@@
$L_{p}^{\prime}(f,k/2)=\displaystyle\int_{{\mathbb{Z}}_{p}^{\times}}g(x)~\!d\mu_{f}(x)$
with $g(x)$ a locally analytic function on ${\mathbb{Z}}_{p}^{\times}$. To understand how to approximate $L_{p}^{\prime}(f,k/2)$, let’s recall how one extends $\mu_{f}$ from a functional on the space of locally polynomial functions of degree at most k2k-2 (as in (3)) to a distribution on all locally analytic functions.
To this end, assume that $n$ is large enough so that $g(x)$ is analytic on $a+p^{n}{\mathbb{Z}}_{p}$ for all $a$. Then decompose ${\mathbb{Z}}_{p}^{\times}$ as a disjoint union of balls of the form $a_{i}+p^{n}{\mathbb{Z}}_{p}$ with each $a_{i}\in{\mathbb{Z}}$ (depending on $n$). Write $\operatorname{TS}_{a_{i}}(g)$ for the Taylor series expansion of $g$ around $x=a_{i}$ truncated to terms of degree at most k2k-2. One defines
$\mu_{f}(g)$ $:=\lim_{n\to\infty}\sum_{i}\int_{a_{i}+p^{n}{\mathbb{Z}}_{p}}\operatorname{TS}_{a_{i}}(g)~\!d\mu_{f}(x)$
@@
which by Lemma 3.2 and the above discussion converges to $L_{p}^{\prime}(f,k/2)$ as $n\to\infty$. Since $L_{n}$ is defined by values of $\mu_{f}$ on polynomials of degree at most k2k-2, we see from Section 2.2 that $L_{n}$ is a linear combination of period of integrals of $f$, and thus each $L_{n}$ can be computed from the modular symbol attached to $f$.
@@
$L^{\prime}_{p}(f,k/2)-L_{n}$ $=\int_{{\mathbb{Z}}_{p}^{\times}}x^{\frac{k}{2}-1}\log_{p}\langle x\rangle~\!d\mu_{f}(x)-\sum_{\begin{subarray}{c}a=1\\ p\nmid a\end{subarray}}^{p^{n}-1}\mu_{f}(\operatorname{TS}_{a}(x^{\frac{k}{2}-1}\cdot\log_{p}\langle x\rangle),a,p^{n})$
@@
$=\sum_{\begin{subarray}{c}a=1\\ p\nmid a\end{subarray}}^{p^{n}-1}\sum_{m=k-1}^{\infty}c_{m,a}\int_{a+p^{n}{\mathbb{Z}}_{p}}(x-a)^{m}~\!d\mu_{f}(x).$
Thus, using Lemma 2.1 to estimate the integrals $\int(x-a)^{m}~\!d\mu_{f}(x)$ and Lemma 3.3 to estimate the $c_{m,a}$, we see
@@ 4. Summary of algorithm to calculate $\mathcal{L}$}-invariants
    The eigenvalues of $U_{p}$ acting on $A$ must all be of the form $\pm p^{\frac{k}{2}-1}$ with a fixed sign. We calculate the $p$-th Hecke polynomial and use it to determine the sign of $a_{p}=a_{p}(A)$. Since $A$ is new of level $\Gamma_{0}$, the sign $w=w(A)$ of the functional equation can extracted from the sign of the eigenvalue for the Atkin–Lehner operator $w_{Np}$ on $A$. Having determined these two signs, we find a quadratic discriminant $D$ as in Proposition 2.4. Write $\chi=\chi_{D}$ for the quadratic character of conductor $D$.
@@
    is equal to $v_{\mathfrak{p}}({\mathcal{L}})$ for $K_{\mathfrak{p}}:{\mathbb{Q}}_{p}]-$many eigenforms occurring in $A$. So, we store this value, with multiplicity $K_{\mathfrak{p}}:{\mathbb{Q}}_{p}]$. If (8) fails for some $\mathfrak{p}$, then return to step (iii)(b) to compute at a deeper level, eventually replacing $L_{1}$ by $L_{2}$, and so on.
@@ 5.1. Initial plots in tame level one
  • The size of a scatter point represents the datum’s multiplicity. For instance if $v_{p}({\mathcal{L}})=-8$ occurs twice in weight $k$ and $v_{p}({\mathcal{L}})=-10$ occurs four times, then the scatter point at (k,10)(k,-10) is a magnitude larger than the scatter point at (k,8)(k,-8).
@@
Figure 5. 33-adic slopes of $\mathcal{L}$}-invariant slopes in level $\Gamma_{0}(3)$ with respect to weights $6\leq k\leq 508$. (Total data: 10752 eigenforms.)
Figure 6. 55-adic slopes of $\mathcal{L}$}-invariant slopes in level $\Gamma_{0}(5)$ with respect to weights $10\leq k\leq 472$. (Total data: 18560 eigenforms.)
Figure 7. 77-adic slopes of $\mathcal{L}$}-invariant slopes in level $\Gamma_{0}(7)$ with respect to weights $4\leq k\leq 290$. (Total data: 10511 eigenforms.)
@@ 5.3. The sign of $a_{p}$
These split as evenly as they could among the two signs: the 9-9 occurs once for each sign and the 4-4 occurs once for each sign. The 2-2 happens to occur for an eigenform with $a_{3}<0$. For a second example, in weight $k=28$ and level $\Gamma_{0}(3)$, there are an even number of eigenforms and their $\mathcal{L}$}-invariants have slopes
@@
Both plots support our two principles above. For instance, in level $\Gamma_{0}(2)$ the most negative sign dependent $\mathcal{L}$}-invariant slope we found was 7-7. This is extremely close to zero, once we note the most negative $\mathcal{L}$}-invariant slope over the same data range is less than 120-120 (see Figure 3). It seems plausible to expect sign dependent slopes $v$ to satisfy $v\geq-c\log(k)$ where $c$ is a positive constant. (Depending on what? Perhaps just $p$.) It also seems reasonable that the total number of sign dependent slopes in weights $k\leq T$ is $O(T\log(T))$, which is a magnitude smaller than roughly $T^{2}$-many eigenforms of weight $k\leq T$.
@@
$p$ NpNp # weights # data up $k_{\mathrm{mid}}$ % sign dependent up to $k_{\mathrm{mid}}$ # total data % sign dependent
22 22 $391$ $3434$ $4.14$ $13002$ $2.37$
22 66 $165$ $1291$ $9.37$ $4603$ $5.02$
22 $10$ $136$ $1680$ $7.5$ $6264$ $3.64$
22 $14$ $126$ $2245$ $8.06$ $7947$ $3.49$
22 $22$ $103$ $2340$ $4.79$ $9013$ $2.34$
33 33 $252$ $2730$ $4.25$ $10752$ $2.18$
33 66 $164$ $1291$ $6.89$ $4546$ $4.4$
33 $15$ $72$ $962$ $12.47$ $3594$ $6.12$
33 2121 $70$ $1296$ $9.57$ $5040$ $4.88$
33 $33$ $42$ $735$ $14.83$ $2941$ $7.17$
55 55 $243$ $5371$ $3.18$ $20341$ $1.61$
55 $10$ $135$ $1680$ $5.24$ $6173$ $2.64$
55 $15$ $72$ $962$ $8.94$ $3594$ $5.29$
55 $35$ $41$ $883$ $10.31$ $3363$ $6.87$
55 $70$ 2020 $280$ $9.29$ $960$ $5.62$
77 77 $144$ $2664$ $3.38$ $10512$ $1.77$
77 $14$ $90$ $1301$ $7.76$ $4085$ $4.38$
77 2121 $70$ $1296$ $6.64$ $5040$ $3.37$
77 $35$ $41$ $883$ $10.76$ $3363$ $5.68$
77 $70$ 1919 $280$ $10.0$ $874$ $6.41$
@@ 5.4. Galois representations
Figure 13 offers a second plot to visualize how $\overline{\rho}$-filtered data sits within an unfiltered dataset. Namely, let $\overline{\rho}$ be the mod $11$ representation associated with the elliptic curve $X_{0}(11)$. Then we consider all eigenforms $f$ for which $\overline{\rho}_{f}\simeq\overline{\rho}\otimes\omega^{j}$ for some $j=0,1,\dotsc,10$. We call these eigenforms “twists of $X_{0}(11)$ modulo 1111”. Each such $\overline{\rho}_{f}$ is surjective, which makes the context slightly different than the Eisenstein examples above.
@@ 6.1. $\mathcal{L}$}-invariants and local Galois representations
Fix $k\geq 2$, a sign $\pm$, and ${\mathcal{L}}\in\overline{{{\mathbb{Q}}}}_{p}$. We write $D=\overline{{{\mathbb{Q}}}}_{p}e_{1}\oplus\overline{{{\mathbb{Q}}}}_{p}e_{2}$ and define $\overline{{{\mathbb{Q}}}}_{p}$-linear operators $\varphi$ and $N$ on $D$ by
$\varphi$ $=\begin{pmatrix}\pm p^{\frac{k}{2}}&0\\ 0&\pm p^{\frac{k}{2}-1}\end{pmatrix}$
@@
In this way, $D$ is a two-dimensional $\overline{{{\mathbb{Q}}}}_{p}$-linear weakly-admissible filtered $(\varphi,N)$-module in the sense of Fontaine’s $p$-adic Hodge theory of local Galois representations [39]. A theorem of Colmez and Fontaine ([33, Théormème A]) implies there exists a two-dimensional local representation
@@
that is semi-stable and non-crystalline, since $N\neq 0$ on $D$, for which $D_{\mathrm{st}}(r)=D$. The Hodge–Tate weights of $r$ equal $\{0,k-1\}$ and $\det(r)=\omega^{k-1}$. We write $r_{k,{\mathcal{L}}}^{\pm}$ for this representation, or we write $V_{k,{\mathcal{L}}}^{\pm}$ when it is more convenient to reference the vector space upon which the Galois groups acts. Any two-dimensional, semi-stable, and non-crystalline, $\overline{{{\mathbb{Q}}}}_{p}$-linear $\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})$-representation with distinct Hodge–Tate weights is isomorphic to a twist of some such $r_{k,{\mathcal{L}}}^{\pm}$. The representations $r_{k,{\mathcal{L}}}^{\pm}$ are irreducible if and only if $k>2$. See [17, Exemple 3.1.2.2] for a similar summary and references.
@@
To avoid complications below, it is convenient here to include a definition of $r_{k,{\mathcal{L}}}^{\pm}$ for $\mathcal{L}$$\infty$. Namely, $r=r_{k,\infty}^{\pm}$ is the unique two-dimensional, irreducible, crystalline representation of $\operatorname{Gal}(\overline{{{\mathbb{Q}}}}_{p}/{\mathbb{Q}}_{p})$ such that $D_{\mathrm{crys}}(r)=\overline{{{\mathbb{Q}}}}_{p}e_{1}\oplus\overline{{{\mathbb{Q}}}}_{p}e_{2}$ is the weakly-admissible filtered $(\varphi,N)$-module, with $N=0$, and for which